SPECTROSCOPY ON EXTREMELY WEAKLY-BOUND MOLECULES: COOPER PAIRS AND EFIMOV TRIMERS

\underline{CHENG CHIN}, The James Franck institute and the Department of Physics, University of Chicago, Chicago, IL, 60637.

Recent research on ultracold atoms and molecules entered an exciting era in which atomic and molecular interactions can be fully controlled and engineered to simulate a broad range of few- and many-body phenomena in condensed matter, molecular physics and even nuclear physics. In this talk I will present two examples. The first example is degenerate Fermi gas with resonant interactions. Our studies of Fermion pairs from precision radiofrequency spectroscopy provide new views on many-body pairing and phase transitions in strongly interacting Fermi systems^{a}. This work was ranked one of the top 10 science breakthroughs in year 2004 by Science magazine. The second one is the observation of a three-body Efimov state, which was conjectured in 1970 by a Russian physicist V. Efimov in the context of nuclear physics. Efimov trimers exist in general Bose systems when the binary interaction is resonantly enhanced. The observation of an Efimov state^{b} opens up new horizon to explore few-body physics in the quantum degeneracy regime^{c}.

aC. Chin, M. Bartenstein, A. Altmeyer, S. Riedl, S. Jochim, J. Hecker Denschlag, and R. Grimm, Science \underline{305}, 1128 (2004)

bT. Kraemer, M. Mark, P. Waldburger, J. G. Danzl, C. Chin, B. Engeser, A. D. Lange, K. Pilch, A. Jaakkola, H.-C. Nägerl and R. Grimm, Nature \underline{440}, 315 (2006)

cC. Chin, T. Kraemer, M. Mark, J. Herbig, P. Waldburger, H.-C. Nägerl, and R. Grimm, Phys. Rev. Lett. \underline{94}, 123201 (2005).